OpenStudy (anonymous):

The demand function and the cost function for x units of a product are p=30/sqrt(x) and c=.95x+4750 Find value of the marginal profit when x=8100. The value of the marginal profit when x= 8100 is

OpenStudy (anonymous):

"demand function and the cost function for units of a product are ...and ,,," !?!

OpenStudy (anonymous):

i edited it

OpenStudy (anonymous):

Do you know how to find Profit?

OpenStudy (anonymous):

not really im really confused. if you could show me step by step itll be helpful

OpenStudy (anonymous):

Big big Hint: Relation between Revenue and Cost

OpenStudy (anonymous):

revnue-cost

OpenStudy (anonymous):

Yep, now what's Revenue?

OpenStudy (anonymous):

30/sqrt(x)?

OpenStudy (anonymous):

nevermind im not sure

OpenStudy (anonymous):

pretty big Hint: Relation between demand and unit!

OpenStudy (cruffo):

for the demand function: input variable x is quantity, the output variable p is price?

OpenStudy (anonymous):

-> R(x) = P (x) * x

OpenStudy (cruffo):

p=30/sqrt(x) means price = 30/sqrt(quantity)

OpenStudy (anonymous):

so 30/sqrt(8100)* 8100?

OpenStudy (anonymous):

Can you find the Profit function now?

OpenStudy (anonymous):

is it the equation i said?

OpenStudy (anonymous):

P (x) = R(x) - C(x)

OpenStudy (anonymous):

so p(x)= 30/sqrt(8100)-(.95(8100) +4750?

OpenStudy (anonymous):

Do NOT plug the numbers in when you don't understand the concept !!!

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so what do i do?

OpenStudy (anonymous):

P (x) = R(x) - C(x) What do you understand about my explanation?

OpenStudy (anonymous):

do you have to find the derivative?

OpenStudy (anonymous):

R(x) = ...?

OpenStudy (anonymous):

p(x)*x

OpenStudy (anonymous):

Yes, now replace the function in!

OpenStudy (anonymous):

30/sqrt(x)* X?

OpenStudy (anonymous):

Wrong multiply: R(x) : 30x/√x = ....?

OpenStudy (anonymous):

do the x cancel out?

OpenStudy (anonymous):

Note: x = √x * √x

OpenStudy (anonymous):

so whats next?

OpenStudy (anonymous):

So how do you cancel?

OpenStudy (anonymous):

multiple the top and bottom by sqrt(x)?

OpenStudy (anonymous):

R(x) : 30x/√x =30√x * √x / √x

OpenStudy (anonymous):

and what do we do next

OpenStudy (anonymous):

What do you cancel out?

OpenStudy (anonymous):

im planing on going to help with the professor by monday, but my assignment is due tonight at 11

OpenStudy (anonymous):

would it be 30sqrt(x)

OpenStudy (anonymous):

Yes, ( finally !!! )

OpenStudy (anonymous):

Now, plug the value in P (x) = R(x) - C(x)

OpenStudy (anonymous):

30sqrt(x)-.95x+4750?

OpenStudy (anonymous):

P (x) = R(x) - C(x) =30√x -( .95x+4750 )

OpenStudy (anonymous):

Do you see your mistake?

OpenStudy (anonymous):

i had to put peranthesis

OpenStudy (anonymous):

=30√x - ( .95x+4750 ) =30√x - .95x - 4750

OpenStudy (anonymous):

the negative sign

OpenStudy (anonymous):

i understand

OpenStudy (anonymous):

Yes, you need to learn writing in equation.

OpenStudy (anonymous):

Now, how to find Marginal Profit?

OpenStudy (anonymous):

im not sure lol

OpenStudy (anonymous):

Why you don't know anything Profit, Revenue, Marginal ....?!!!

OpenStudy (anonymous):

Marginal Profit P'(x)

OpenStudy (anonymous):

i was sick so i missed the lessons, but the assignment is due tonight at 11pm which is on the computer and have to hand in on time. meeting with the professor on monday

OpenStudy (anonymous):

(15/sqrt(x)) + 95?

OpenStudy (anonymous):

This is not your fist post about finding Marginal, and many of solvers have already helped you out!

OpenStudy (anonymous):

Take derivative of P (x) -> P' (x) = ....?

OpenStudy (anonymous):

(15/sqrt(x))+15?

OpenStudy (anonymous):

+95*

OpenStudy (anonymous):

P(x) = 30√x - .95x -> P'(x) = 15/√x - .95 Thus P'(8100) = ....

OpenStudy (anonymous):

-47/60?

OpenStudy (anonymous):

i got it. thanks for the help